OpenStudy (anonymous):

Is the k in Coulomb's law F = k qq'/r2 same as K = ε_0/ε_r ?

OpenStudy (anonymous):

No. In SI units, the k in Coulombs law is \[ k = \frac{1}{4\pi \epsilon_0} \approx 8.99\times 10^9 \space Nm^2/C^2\]

OpenStudy (anonymous):

What is it called as ? Is is called as permeability of the medium?

OpenStudy (anonymous):

No, it's called the Coulomb constant.

OpenStudy (anonymous):

It's often denoted by \[k_c \text{ or } k_e\] if there's a chance for ambiguity.

OpenStudy (anonymous):

What is called as permeability of a medium and permittivity of a medium?

OpenStudy (anonymous):

The permittivity of a medium is \[ \epsilon \] and the permeability of a medium is \[\mu\] If you're talking about vacuum, they are denoted \[\epsilon_0 \text{ and } \mu_0 \] respectively.

OpenStudy (anonymous):